Given that C₁=10.0F, C₂=5.0F, and C3=4.0F, find the total capacitance for each connection shown.


1
Expert's answer
2022-03-30T13:51:20-0400

Parrallel combination



"C_{eq}=C_1+C_2+C_3=(10+5+4)\mu F=19\mu F"

Series combination

"\frac{1}{C_{eq}}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C_{eq}}=\frac{1}{10}+\frac{1}{5}+\frac{1}{4}=\frac{10+20+25}{100}"

"C_{eq}=\frac{100}{55} \mu F"

"C_{eq}=1.81\mu F"